2x+y=4 I am trying to find the x and y intercept of the equation

Answer by whty89(20)

To find the x and y intercepts plug in 0 for the other variable and solve the equation
2x+y=4
2(0)+y=4 2x+(0)=4
y=4 x=2
x intercept is (2,0)
y intercept is (0,4)